OpenStudy (anonymous):

Choose the equation that could be used to find two consecutive integers whose sum is 69. 1) n+(n+2)=69 2) n+(n+1)=69 3) n-(n+1)=69 4) n+2n= 69 Could really use the help guys! Thank you :)

OpenStudy (er.mohd.amir):

n+n+1=69 2n=69-1 n=68/2 n=34 so numbers 34 and 35

OpenStudy (anonymous):

Thank you so much! I got it right because of you! :)